23. An object traveling with a speed of 10 m/s slides on a horizontal floor. How far will it travel before coming to rest if the coefficient of friction is 0.30?


1
Expert's answer
2021-09-30T10:21:22-0400

Given:

"v_i=10\:\rm m/s"

"v_f=0\:\rm m/s"

"\mu=0.30"


The acceleration of the object

"a=\frac{F}{m}=\frac{-\mu mg}{m}=-\mu g"

The distance traveled by the object

"d=\frac{v_f^2-v_i^2}{2a}=\frac{v_f^2-v_i^2}{-2\mu g}"

"d=\frac{0^2-10^2}{-2*0.30*9.8}=17\:\rm m"